# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 2.5. Holonomy maps of a submetry between Alexandrov spaces
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces and suppose $\alpha\colon[0,b]\to Y$ is a geodesic such that for each $x\in f^{-1}(\alpha(0))$, its initial direction has a unique lift in $H_{x}$ and the holonomy map $p_{\alpha}$ is well defined. Then the holonomy map $p_{\alpha}$ is continuous and surjective.
@@
We first prove the surjectivity of the holonomy map $p_{\alpha}$. Consider $z\in f^{-1}(\alpha(b))$. Then, for the geodesic $\alpha^{-1}(t)=\alpha(b-t)$, we have a unique horizontal lift $\widetilde{\alpha^{-1}}$ in $X$. We consider $x=\widetilde{\alpha^{-1}}(b)$, and observe that $f(x)=\alpha(0)$ by construction. The geodesic $(\widetilde{\alpha^{-1}})^{-1}(t)=\widetilde{\alpha^{-1}}(b-t)$ is a horizontal lift of $\alpha$ starting at $x$. Since the horizontal lifts of $\alpha$ are unique by hypothesis, we conclude that $p_{\alpha}(x)=\widetilde{\alpha^{-1}}(b-b)=z$.
We now prove that $p_{\alpha}$ is continuous. Fix $x\in f^{-1}(\alpha(0))$. Consider a sequence $\{x_{i}\}_{i\in\mathbb{N}}\subset f^{-1}(\alpha(0))$ converging to $x$ under the induced metric of $f^{-1}(\alpha(0))$. Since the intrinsic metric is Lipschitz equivalent to the induced metric, we have $d_{X}(x_{i},x)\to 0$ as $i\to\infty$. Denote by $\tilde{\alpha}_{i}$ and $\tilde{\alpha}$ the unique horizontal lifts of $\alpha$ passing through $x_{i}$ and $x$, respectively. By construction, for $t\in[0,b]$, we have $d_{X}(x_{i},\tilde{\alpha}_{i}(t))=d_{Y}(\alpha(0),\alpha(t))=d_{X}(\tilde{\alpha}(t),x)$. Thus, by the continuity of the metric,
$t=d_{X}(\tilde{\alpha}(t),x)$ $=d_{Y}(\alpha(0),\alpha(t))$
$=\lim_{i\to\infty}d_{X}(x_{i},\tilde{\alpha}_{i}(t))$
$=d_{X}\left(\lim_{i\to\infty}x_{i},\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)$
$=d_{X}(x,\lim_{i\to\infty}\tilde{\alpha}_{i}(t)).$
This implies that the curve $\beta(t)=\lim_{i\to\infty}\tilde{\alpha}_{i}(t)$ is well defined. By the continuity of $f$,
$(f\circ\beta)(t)=f\left(\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)=\lim_{i\to\infty}f(\tilde{\alpha}_{i}(t))=\lim_{i\to\infty}\alpha(t)=\alpha(t).$
Moreover, for any $s,t\in[0,b]$, we have
$d_{X}(\tilde{\alpha}(t),\tilde{\alpha}(s))$ $=|s-t|$
$=d_{Y}(\alpha(s),\alpha(t))$
$=\lim_{i\to\infty}d_{X}(\tilde{\alpha}_{i}(s),\tilde{\alpha}_{i}(t))$
$=d_{X}\left(\lim_{i\to\infty}\tilde{\alpha}_{i}(s),\lim_{i\to\infty}\tilde{\alpha}_{i}(t)\right)$
$=d_{X}(\beta(s),\beta(t).$
From this we conclude that $\beta$ is a geodesic in $X$ and it is a horizontal lift of $\alpha$ starting at $\beta(0)=\lim_{i\to\infty}\tilde{\alpha}_{i}(0)=\lim_{i\to\infty}x_{i}=x$. By the uniqueness of the horizontal lift of $\alpha$, we conclude that $\tilde{\alpha}=\beta$. In particular,
$\lim_{i\to\infty}p_{\alpha}(x_{i})=\lim_{i\to\infty}\tilde{\alpha}_{i}(b)=\beta(b)=\tilde{\alpha}(b)=p_{\alpha}(x).$
Thus, $p_{\alpha}$ is continuous. ∎
@@
2.6. Isometry groups of Alexandrov spaces
As for Riemannian manifolds, the isometry group of an Alexandrov space admits a Lie group structure. Additionally, there is an optimal upper bound on the dimension of the isometry group, and a rigidity result holds when this bound is attained.
Theorem 2.18 ([17, Theorem 1.1] and [18, Theorems 3.1 and 4.1]).
@@ 2.6. Isometry groups of Alexandrov spaces
The conclusions of Theorem 2.18 (1) also hold for a complete length space $(X,d)$ which is locally an Alexandrov space. That is, for each $x\in X$, there exist a neighborhood $U_{x}\subset X$ and a real number $k_{x}$ such that $(U_{x},{\left.\kern-1.2ptd\right|_{U_{x}}})$ is an Alexandrov space with $\mathrm{curv}\geq k$.
@@
Let $g\colon X\to X$ be an isometry of an Alexandrov space $X$, and suppose that there exists $x\in X$ such that $g(x)=x$. Then the differential $dg_{x}\colon T_{x}X\to T_{x}X$ induces an isometry on the space of directions $\Sigma_{x}(X)\subset T_{x}X=C\Sigma_{x}(X)$. Let $G\subset\operatorname{Isom}(X)$ be a group acting on $X$ isometrically. Given $x\in X$, we define the orbit of $G$ through $x$ as $G(x)=\{g\cdot x\mid g\in G\}$. The isotropy subgroup at $x$ is the group $G_{x}=\{g\in G\mid g\cdot x=x\}$. The isotropy representation of $G_{x}$ into $\operatorname{Isom}(\Sigma_{x}(X))$ is given by setting $g\cdot v=dg_{x}(v)$ for $g\in G_{x}$ and $v\in\Sigma_{x}(X)$. The action of $G$ on $X$ is effective if $\cap_{x\in X}G_{x}=\{e\}$. The following lemma implies that, if $G$ acts effectively on $X$, then, for any $x\in X$, the action of $G_{x}$ on $\Sigma_{x}(X)$ given by the isotropy representation is also effective.
Lemma 2.20 ([18, Lemma 3.2]).
@@
Let $X$ be an Alexandrov space and let $g\colon X\to X$ an isometry. If there is some $x_{0}\in X$ such that $g(x_{0})=x_{0}$ and $dg_{x_{0}}\colon\Sigma_{x_{0}}(X)\to\Sigma_{x_{0}}(X)$ is the identity, then $g(x)=x$ for all $x\in X$.
2.7. Foliated homeomorphisms
@@
To conclude this section, let us recall some results on foliated homeomorphisms (cf. [32, Section 3]). Let $p\colon X\to Y$ be a quotient map between two topological spaces $X$ and $Y$. Denote by =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$ the partition of $X$ induced by the preimages of $p$. A continuous map $h\colon X\to X$ is an $\mathcal{F}$-map if, for some $L\in\mathcal{F}$, we have $h(L)\subset L^{\prime}$ for some $L^{\prime}\in\mathcal{F}$. Thus, every $\mathcal{F}$-map $h\colon X\to X$ induces a well defined map $\psi(h)\colon Y\to Y$ making the following diagram commute:
${X}$${X}$${Y}$${Y}$$\scriptstyle{p}$$\scriptstyle{h}$$\scriptstyle{p}$$\scriptstyle{\psi(h)}$
@@
We denote by $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ the group of all $\mathcal{F}$-homeomorphisms (i.e., all $\mathcal{F}$-maps with an inverse map which is an $\mathcal{F}$-map). We let $\operatorname{\mathrm{Homeo}}(Y)$ be the group of all self-homeomorphisms of $Y$. We have a group homomorphism $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$. The map $p\colon X\to Y$ admits local cross-sections if, for any $y\in Y$, there exists an open neighborhood $V\subset Y$ and a continuous map $\sigma\colon V\to X$ such that $p\circ\sigma=\mathrm{Id}_{V}$. When $Y$ is locally compact and Hausdorff, the following theorem guarantees the continuity of $\psi$ when we equip both $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ and $\operatorname{\mathrm{Homeo}}(Y)$ with the compact-open topology.
Theorem 2.21 ([32, Corollary 3.4]).
@@
Let $X,Y$ be topological spaces with $Y$ locally compact and Hausdorff. If $p\colon X\to Y$ is a quotient map, then each of the following conditions implies that the homomorphism $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ is continuous:
@@
    The map $p$ is open map and admits local cross-sections.
@@
3. Proof of Theorem A
We will prove each item in Theorem A separately. To begin, we prove some auxiliary results on proper metric submetries.
@@ 3. Proof of Theorem A
Let $p\colon X\to Y$ be a proper submetry from a length space $X$ onto a metric space $Y$ and set =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$. If the group $\operatorname{Isom}(Y)$ is closed, then the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries is a closed subset of $\operatorname{Isom}(X)$.
@@
Let $\{f_{n}\}\subset\operatorname{Isom}(X,\mathcal{F})\subset\operatorname{\mathrm{Homeo}}(X)$ be a sequence of foliated isometries of $X$ and suppose that there exists an isometry $f\in\operatorname{Isom}(X)$ such that $f_{n}\to f$ in the compact-open topology of $\operatorname{\mathrm{Homeo}}(X)$. We will show that $f$ is a foliated map.
Each foliated isometry $f_{n}$ of $X$ induces an isometry $f^{\ast}_{n}$ of $Y$, defined as $f^{\ast}_{n}(x^{\ast})=f^{\ast}_{n}(p(x))=p(f_{n}(x))$, where $x\in X$ is any point in the preimage of $x^{\ast}$ under the submetry $p\colon X\to Y$. Since $f_{n}$ is a foliated map, $f^{\ast}_{n}$ is well defined, i.e., its definition is independent of the choice of $x\in p^{-1}(x^{\ast})$.
Consider the map $\psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ mapping a foliated homeomorphism $h\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ to $\psi(h)=p\circ h$. By Theorem 2.21, $\psi$ is continuous with respect to the compact-open topology. Restricting $\psi$ to the foliated isometries yields a map $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ that is also continuous with respect to the compact-open topology.
Fix $x,x^{\prime}\in X$ such that $p(x)=p(x^{\prime})$. By construction, $f_{n}(x)\to f(x)$ and $f_{n}(x^{\prime})\to f(x^{\prime})$ as $n\to\infty$. Since $p$ is continuous we have that $p(f_{n}(x))\to p(f(x))$ and $p(f_{n}(x^{\prime}))\to p(f(x^{\prime}))$ as $n\to\infty$. Given that $f_{n}\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$, then we have that $p(f_{n}(x))=p(f_{n}(x^{\prime}))$. By uniqueness of the limit, we have that $p(f(x))=p(f(x^{\prime}))$. This implies that for $x,x^{\prime}\in X$ such that $p(x)=p(x^{\prime})$ we have $p(f(x))=p(f(x^{\prime}))$, and thus $f\in\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$. Thus, $f$ is a foliated isometry. ∎

### v3

@@ 2.5 Holonomy maps of a submetry between Alexandrov spaces
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces and suppose $\alpha\colon[0,b]\to Y$ is a geodesic such that for each $x\in f^{-1}(\alpha(0))$ its initial direction has a unique lift in $H_{x}$, and the holonomy map $p_{\alpha}$ is well defined. Then the holonomy map $p_{\alpha}$ is continuous and surjective.
@@
We first prove the surjectivity of the holonomy map $p_{\alpha}$. Consider $z\in f^{-1}(\alpha(b))$. For the geodesic $\alpha^{-1}(t)=\alpha(b-t)$, there exists a horizontal lift $\widetilde{\alpha^{-1}}$ in $X$ with $\widetilde{\alpha^{-1}}(0)=z$. Set $x=\widetilde{\alpha^{-1}}(b)$, and observe that $f(x)=\alpha(0)$ by construction. The geodesic $\smash{\bigl(\widetilde{\alpha^{-1}}\bigr)^{-1}(t)}=\smash{\widetilde{\alpha^{-1}}(b-t)}$ is a horizontal lift of $\alpha$ starting at $x$. Since the horizontal lifts of $\alpha$ are unique by hypothesis, we conclude that $p_{\alpha}(x)=\widetilde{\alpha^{-1}}(0)=z$.
We now prove that $p_{\alpha}$ is continuous. Fix $x\in f^{-1}(\alpha(0))$ and let $\{x_{i}\}_{i\in\mathbb{N}}\subset f^{-1}(\alpha(0))$ be a sequence converging to $x$ in the intrinsic metric of the fiber. By Theorem 2.3, the intrinsic and the induced metric are locally Lipschitz equivalent on the fiber, hence also $d_{X}(x_{i},x)\to 0$ as $i\to\infty$}. Let $\tilde{\alpha}_{i}\colon[0,b]\to X$ be the unique horizontal lift of $\alpha$ with $\tilde{\alpha}_{i}(0)=x_{i}$. Let $\tilde{\alpha}\colon[0,b]\to X$ be the unique lift of $\alpha$ with $\tilde{\alpha}(0)=x$. To simplify the notation, set $y_{i}=p_{\alpha}(x_{i})=\tilde{\alpha}_{i}(b)$ and $y=p_{\alpha}(x)=\tilde{\alpha}(b)$}. We will show that $y_{i}\to y$.
For each $t\in[0,b]$,
$d_{X}(\tilde{\alpha}_{i}(t),x)\leq d_{X}(\tilde{\alpha}_{i}(t),x_{i})+d_{X}(x_{i},x)\leq b+d_{X}(x,x_{i}).$
Thus, for all sufficiently large $i$, we have that $\tilde{\alpha}_{i}([0,b])$ is contained in the closed ball $B_{b+1}(x)$, which is a compact subset of $X$. In particular, $\{y_{i}\}_{i\in\mathbb{N}}$ is contained in a compact subset of $X$.
Let $\{y_{i_{j}}\}_{i\in\mathbb{N}}$ be an arbitrary convergent subsequence with $y_{i_{j}}\to y^{\prime}$ in $X$. Let us show that $y^{\prime}=y$}. Since the geodesics $\tilde{\alpha}_{i_{j}}\colon[0,b]\to B_{b+1}(x)\subset X$ have the same length and are contained in a compact subset of $X$, the Arzelà–Ascoli theorem (see [11, Theorem 2.5.14]) implies that there exists a sub-subsequence (which we will not relabel) $\tilde{\alpha}_{i_{j}}$ that converges uniformly to a curve $\beta\colon[0,b]\to X$, which must necessarily be a geodesic (see [11, Proposition 2.5.17]). By the continuity of $f$,
$f\circ\beta=\lim_{j\to\infty}f\circ\tilde{\alpha}_{i_{j}}=\lim_{j\to\infty}\alpha=\alpha.$
Hence, $\beta$ is a horizontal lift of $\alpha$. Observe now that
$\beta(0)=\lim_{j\to\infty}\tilde{\alpha}_{i_{j}}(0)=\lim_{j\to\infty}x_{i_{j}}=x.$
Since we have assumed that lifts starting at the same point are unique, we must have $\beta=\tilde{\alpha}$. Taking endpoints, we get
$y^{\prime}=\lim_{j\to\infty}y_{i_{j}}=\lim_{j\to\infty}\tilde{\alpha}_{i_{j}}(b)=\beta(b)=\tilde{\alpha}(b)=y.$
Therefore, every convergent subsequence of $\{y_{i}\}_{i\in\mathbb{N}}$ converges to $y$. Since $\{y_{i}\}_{i\in\mathbb{N}}$ is contained in a compact subset of $X$, it follows that $y_{i}\to y$, i.e., $p_{\alpha}(x_{i})\to p_{\alpha}(x)$. Thus, $p_{\alpha}$ is continuous. ∎
@@
2.6 Isometry groups of Alexandrov spaces
As for Riemannian manifolds, the isometry group of an Alexandrov space admits a Lie group structure. Additionally, there is an optimal upper bound on the dimension of the isometry group and a rigidity result holds when this bound is attained.
Theorem 2.18 ([22, Theorem 1.1] and [23, Theorems 3.1 and 4.1]).
@@ 3. $(3)$
The conclusions of Theorem 2.18(1)(1) also hold for a complete length space $(X,d)$ which is locally an Alexandrov space. That is, for each $x\in X$, there exist a neighborhood $U_{x}\subset X$} and a real number $k_{x}$ such that $(U_{x},{\left.\kern-1.2ptd\right|_{U_{x}}})$ is an Alexandrov space with $\mathrm{curv}\geq k$.
@@
Let $g\colon X\to X$ be an isometry of an Alexandrov space $X$, and suppose that there exists $x\in X$} such that $g(x)=x$. Then the differential ${\rm d}g_{x}\colon T_{x}X\to T_{x}X$ induces an isometry on the space of directions $\Sigma_{x}(X)\subset T_{x}X=C\Sigma_{x}(X)$. Let $G\subset\operatorname{Isom}(X)$ be a group acting on $X$ isometrically. Given $x\in X$, we define the orbit of $G$ through $x$ as $G(x)=\{g\cdot x\mid g\in G\}$. The isotropy subgroup at $x$ is the group $G_{x}=\{g\in G\mid g\cdot x=x\}$. The isotropy representation of $G_{x}$ into $\operatorname{Isom}(\Sigma_{x}(X))$ is given by setting $g\cdot v={\rm d}g_{x}(v)$ for $g\in G_{x}$ and $v\in\Sigma_{x}(X)$. The action of $G$ on $X$ is effective if $\bigcap_{x\in X}G_{x}=\{e\}$. The following lemma implies that, if $G$ acts effectively on $X$, then for any $x\in X$} the action of $G_{x}$ on $\Sigma_{x}(X)$ given by the isotropy representation is also effective.
Lemma 2.20 ([23, Lemma 3.2]).
@@
Let $X$ be an Alexandrov space and let $g\colon X\to X$ an isometry. If there is some $x_{0}\in X$ such that $g(x_{0})=x_{0}$ and ${\rm d}g_{x_{0}}\colon\Sigma_{x_{0}}(X)\to\Sigma_{x_{0}}(X)$ is the identity, then g(x)=x{g(x)=x} for all $x\in X$.
Let $Z$ be a connected locally compact metric space (in particular, any Alexandrov space). The action of $\operatorname{Isom}(Z)$ on $Z$ is proper when $\operatorname{Isom}(Z)$ is equipped with the topology of pointwise convergence (see [42, Proposition in Section 4, p. 11]). This topology agrees with the compact-open topology, which in turn is equivalent to the topology of uniform convergence over compact subsets (see, for example, [41]). Hence, the action is proper for the compact-open topology as well. We record this fact for use in the proof of Theorem B, and note that one may also prove it for Alexandrov spaces by adapting the Riemannian proof (see, for example, [1, Proposition 3.62] and cf. [18]).
Proposition 2.21.
Let $X$ be a connected Alexandrov space. Then the action of $\operatorname{Isom}(X)$ ((with the compact-open topology)) on $X$ is proper.
2.7 Foliated maps
@@
To conclude this section, let us recall some results on foliated homeomorphisms (cf. [40, Section 3]). Let $p\colon X\to Y$ be a quotient map between two topological spaces $X$ and $Y$. Denote by $\mathcal{F}=\big\{p^{-1}(y)\mid y\in Y\big\}$} the partition of $X$ induced by the pointwise preimages of $p$. A continuous map $h\colon X\to X$ is foliated if, for any $L\in\mathcal{F}$, we have $h(L)\subset L^{\prime}$ for some $L^{\prime}\in\mathcal{F}$. Thus, every foliated map $h\colon X\to X$ induces a well-defined map $\Psi(h)\colon Y\to Y$ given by $\Psi(h)(p(x))=p(h(x))$ making the following diagram commute
[…]
@@
We denote by $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ the group of all foliated homeomorphisms (i.e., all foliated maps with an inverse map which is also foliated). We let $\operatorname{\mathrm{Homeo}}(Y)$ be the group of all self-homeomorphisms of $Y$. We have a group homomorphism $\Psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ given by $f\mapsto\Psi(f)$. The map $p\colon X\to Y$ admits local cross-sections if, for any $y\in Y$, there exists an open neighborhood $V\subset Y$ and a continuous map $\sigma\colon V\to X$ such that $p\circ\sigma=\mathrm{Id}_{V}$. When $Y$ is locally compact and Hausdorff, the following theorem guarantees the continuity of $\Psi$ when we equip both $\operatorname{\mathrm{Homeo}}(X,\mathcal{F})$ and $\operatorname{\mathrm{Homeo}}(Y)$ with the compact-open topology.
Theorem 2.22 ([40, Corollary 3.4]).
@@
Let $X$ and $Y$ be topological spaces with $Y$ locally compact and Hausdorff. If $p\colon X\to Y$ is a quotient map, then each of the following conditions implies that the homomorphism $\Psi\colon\operatorname{\mathrm{Homeo}}(X,\mathcal{F})\to\operatorname{\mathrm{Homeo}}(Y)$ is continuous:
@@ 2. $(2)$
    The map $p$ is open and admits local cross-sections.
@@ 3. $(3)$
Suppose now that $X$ is a metric space, the fibers of the quotient map $p\colon X\to Y$ are closed and equidistant (i.e., $d_{L_{i}}$ is constant on $L_{j}$, for $i,j=1,2$), and the quotient space $Y=X/\mathcal{F}$ is equipped with the quotient metric defined by
$d_{Y}(p(L_{1}),p(L_{2}))=\inf\{d_{X}(x_{1},x_{2})\mid x_{i}\in L_{i}\}$
for fibers $L_{1},L_{2}\in\mathcal{F}$. Denote by $\operatorname{Isom}(X,\mathcal{F})$ the group of foliated isometries, i.e., isometries of $X$ sending leaves to leaves. Then the map $\Psi(f)\in\operatorname{\mathrm{Homeo}}(Y)$ is an isometry of $Y$. Indeed, let $y_{i}=p(L_{i})$} with leaves $L_{i}\in\mathcal{F}$, and fix $x_{i}\in L_{i}$, $i=1,2$. Then
$d_{Y}\bigl(\Psi(f)(y_{1}),\Psi(f)(y_{2})\bigr)$ $=d_{Y}\bigl(p(f(x_{1})),p(f(x_{2}))\bigr)=d_{X}\bigl(f(L_{1}),f(L_{2})\bigr)$
$=d_{X}(L_{1},L_{2})=d_{Y}\bigl(p(L_{1}),p(L_{2})\bigr)=d_{Y}(y_{1},y_{2}).$
Thus, $\Psi(f)\in\operatorname{Isom}(Y)$. Consequently, $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ is a group homomorphism.
We conclude this section with the following general facts, which we will use in the proofs of Theorems A and B.
@@
Let $p\colon X\to Y$ be a continuous surjective map between metric spaces and set $\mathcal{F}=\big\{p^{-1}(y)\mid y\in Y\big\}$}. Then the group $\operatorname{Isom}(X,\mathcal{F})$ of foliated isometries is closed in $\operatorname{Isom}(X)$.
@@
Recall that $\operatorname{Isom}(X)$ is equipped with the compact-open topology. For each ordered pair $(x,x^{\prime})\in X\times X$ with $p(x)=p(x^{\prime})$, define
$\Phi_{x,x^{\prime}}\colon\ \operatorname{Isom}(X)\to Y\times Y,\qquad g\mapsto\bigl(p(gx),\,p(gx^{\prime})\bigr).$
In the compact-open topology, for each point $x\in X$, the Myers–Steenrod map (i.e., the evaluation map)
$\mu_{x}\colon\ \operatorname{Isom}(X)\to X,\qquad g\mapsto gx$
is continuous (cf. [23, 29]). Therefore, the map $\Phi_{x,x^{\prime}}=(p\circ\mu_{x})\times(p\circ\mu_{x^{\prime}})$ is continuous.
Since $Y$ is a metric space (hence Hausdorff), the diagonal $\Delta_{Y}=\{(y,y)\in Y\times Y\}$ is closed in $Y\times Y$}. An isometry $g\in\operatorname{Isom}(X)$ is foliated if and only if, for all $x,x^{\prime}\in X$, $p(x)=p(x^{\prime})$ implies that $p(gx)=p(gx^{\prime})$, i.e., if and only if $\Phi_{x,x^{\prime}}(g)\in\Delta_{Y}$ for every pair $(x,x^{\prime})$ with $p(x)=p(x^{\prime})$}. Therefore,
$\operatorname{Isom}(X,\mathcal{F})=\bigcap\Phi_{x,x^{\prime}}^{-1}(\Delta_{Y}),$
where the intersection is taken over the set of all pairs $(x,x^{\prime})\in X\times X$ with $p(x)=p(x^{\prime})$. Since $\Delta_{Y}$ is closed and $\Phi_{x,x^{\prime}}$ is continuous, each preimage $\Phi_{x,x^{\prime}}^{-1}(\Delta_{Y})$ is closed in $\operatorname{Isom}(X)$. Hence, $\operatorname{Isom}(X,\mathcal{F})$ is a closed subgroup of $\operatorname{Isom}(X)$. ∎